OpenStudy (anonymous):

Which of the following expresses the circumference of a circle in terms of A, its area?

hartnn (hartnn):

C=2pir,A=pi r^2 just divide.

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

Let me get a pen...

hartnn (hartnn):

ok, square C

hartnn (hartnn):

then put the value of r^2 from A=pi r^2 in your squared expression.,

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

2pir/pir squared?

OpenStudy (anonymous):

how did you do it again?

hartnn (hartnn):

give your options...

OpenStudy (anonymous):

2A,2A sqrt pi, 2pi sqrt A, 2 sqrt pi times A, 2piA

hartnn (hartnn):

your options has r in it ?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it has sqrt

hartnn (hartnn):

nice, take square of C

OpenStudy (anonymous):

square or square root?

hartnn (hartnn):

we will take square root afterwards, after we eliminate r, first take square

OpenStudy (anonymous):

(2pir) quantity squared?

hartnn (hartnn):

yup. find r^2 from A=pi r^2

OpenStudy (anonymous):

I got 4pi when I divided 4pi sq. r sq./pir sq.

OpenStudy (anonymous):

Is that right?

hartnn (hartnn):

didn't get u,but r^2 = A/pi. C^2 = 4 pi^2 r^2 =4 pi^2 A/pi got this ?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

you also sq. the circumference

OpenStudy (anonymous):

so hold up

OpenStudy (anonymous):

C sq.=4 pi sq. r sq.?

OpenStudy (anonymous):

right?

hartnn (hartnn):

yup

OpenStudy (anonymous):

Where did you get A/pi?

hartnn (hartnn):

A = pi r^2

OpenStudy (anonymous):

did you divided circumference sq. by area?

hartnn (hartnn):

i found r^2 from A and put it in C^2

OpenStudy (anonymous):

cause I got 4pi when I did that

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so you solved for radius sq.?

hartnn (hartnn):

yup.

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

I am still confused

OpenStudy (anonymous):

how did we solve for r squared?

hartnn (hartnn):

ok. then what are u comfortable with?

OpenStudy (anonymous):

I know that we sq. C to get 4pi sq. r sq.

OpenStudy (anonymous):

so what next?

hartnn (hartnn):

A=pi r^2 so whats r^2 from here ?

OpenStudy (anonymous):

A/pi?

OpenStudy (anonymous):

A/pi= r sq.?

hartnn (hartnn):

yup, now put that in C sq.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

4 pi sq. times A/pi?

hartnn (hartnn):

now cancel 1 pi

OpenStudy (anonymous):

4 pi times A?

hartnn (hartnn):

yes. that is C^2 so what is C ?

OpenStudy (anonymous):

2 times sq. pi times A?

hartnn (hartnn):

2 sqrt pi times A yup

OpenStudy (anonymous):

phew, that problem took a while